OpenStudy (anonymous):

how do I solve this the product of 3 less than twice x and 2 more than the quantity 3 times x ?

OpenStudy (anonymous):

i cant seem to figure this problem

OpenStudy (anonymous):

i got 1/5 what did others get? i dont want to explain it if im wrong

OpenStudy (anonymous):

It's a product, so you first need to figure out what and what you're multiplying : 3 less than twice x * 2 more than 3 times x So far so good? @imasofty I don't think this will give you a number at the end, unless I'm mistaken. That should give you a polynomial expression

OpenStudy (anonymous):

oh i just set them equal and solved for x

OpenStudy (anonymous):

oh snap its a product so you aregoing to be foiling have fun with that

OpenStudy (anonymous):

To nudge you in the right direction, here's how to "mathematify" the first term : 3 less than twice x. twice x is 2x. 3 less than 2x is 2x-3 So we now have this : (2x-3) * (2 more than 3 times x) Think you can give that second term a try?

OpenStudy (anonymous):

3x-2

OpenStudy (anonymous):

Im probably wrong

OpenStudy (anonymous):

close, but not quite. 3 times x is 3x, that part you got. 2 more than 3x will be what?

OpenStudy (anonymous):

6x

OpenStudy (mertsj):

Two MORE

OpenStudy (anonymous):

8x

OpenStudy (mertsj):

You have 3 dollars and I have 2 more dollars than you. How much do I have and how did you get it?

OpenStudy (anonymous):

if I ask you what 2 more than 5 is, what will you answer?

OpenStudy (anonymous):

5x

OpenStudy (anonymous):

7x

OpenStudy (mertsj):

5 dollars!!! How do you get 5 dollars?

OpenStudy (mertsj):

You make something more than something else by adding to the first quantity.

OpenStudy (mertsj):

You have 3 dollars and I have 2 dollars more 3+2=5. I have 5 dollars

OpenStudy (mertsj):

Similarly 2 more than 3x is 3x+2

OpenStudy (anonymous):

so i'm adding 3x+2=

OpenStudy (mertsj):

Bingo

OpenStudy (anonymous):

alright problem solved

OpenStudy (mertsj):

Three less than twice x is 2x-3 2 more than 3x is 3x+2 Multiply those two binomials because in order to find a product you have to multiply.

OpenStudy (mertsj):

(2x-3)(3x+2)= Use FOIL

OpenStudy (anonymous):

im confused with the inner and outer process